OpenStudy (anonymous):

5/8(8x+32)+1=1/2(12x-2)+15

OpenStudy (anonymous):

x = 7

OpenStudy (anonymous):

5x + 20 + 1 = 6x - 1 + 15 x = 21+ 1 - 15 x = 7